### v1

# Diameter Bounds for Friends-and-Strangers Graphs

Abstract.
Consider two $n$-vertex graphs $X$ and $Y$, where we interpret $X$ as a social network with edges representing friendships and $Y$ as a movement graph with edges representing adjacent positions. The friends-and-strangers graph $\mathsf{FS}(X,Y)$ is a graph on the $n!$ permutations $V(X)\to V(Y)$, where two configurations are adjacent if and only if one can be obtained from the other by swapping two friends located on adjacent positions.
Friends-and-strangers graphs were first introduced by Defant and Kravitz [6], and generalize sliding puzzles as well as token swapping problems. Previous work has largely focused on their connectivity properties. In this paper, we study the diameter of the connected components of $\mathsf{FS}(X,Y)$. Our main result shows that when the underlying friendship graph is a star with $n$ vertices, the friends-and-strangers graph has components of diameter $O(n^{4})$. This implies, in particular, that sliding puzzles are always solvable in polynomially many moves. Our work also provides explicit efficient algorithms for finding these solutions.
We then extend our results to general graphs in two ways. First, we show that the diameter is polynomially bounded when both the friendship and the movement graphs have large minimum degree. Second, when both the underlying graphs $X$ and $Y$ are Erdős-Rényi random graphs, we show that the distance between any pair of configurations is almost always polynomially bounded under certain conditions on the edge probabilities.
1. Introduction
Let $X$ be a “friendship graph” on $n$ vertices, where the vertices represent people and the edges represent friendships. Let $Y$ be a “movement graph” on $n$ locations. Suppose that the $n$ people in $X$ are placed on the vertex set of $Y$, each in a distinct location. Two people can swap positions in a “friendly swap” if they are friends in $X$ and are in adjacent positions in $Y$. The following two problems naturally come to mind.
Question 1.1.
Given two configurations of people on the movement graph, can one be obtained from the other via a sequence of friendly swaps?
Question 1.2.
What is the maximum number of moves needed to go from one configuration to another?
More formally, let $G$ be a graph on $n!$ vertices. We say that $G$ is the friends-and-strangers graph of $X$ and $Y$, denoted as $\mathsf{FS}(X,Y)$, if and only if it satisfies the following conditions:
  1. (1)
    there exists a bijection $f$ from the vertex set of $G$ to the set of permutations mapping the people in $X$ onto $Y$;
  2. (2)
    let $v$ and $w$ be any two vertices in $G$. Then, $v$ and $w$ are adjacent in $G$ if and only if $f(v)$ is reachable from $f(w)$ in exactly one friendly swap.
[…]
More generally, sliding puzzles can be defined on arbitrary $n$-vertex graphs $Y$ by placing n1n-1 tiles on distinct vertices, with one vertex left unoccupied. A single move consists of moving a tile from its current location to an adjacent unoccupied location. The movement graph over configurations of the tiles is then given by $\mathsf{FS}(\mathsf{Star}_{n},Y)$ where the tiles correspond to the leaves of the star graph and the center of the star corresponds to the “hole” representing the unoccupied position. Solving the puzzle corresponds to reaching a target configuration through a sequence of legal moves. Connectivity in $\mathsf{FS}(\mathsf{Star}_{n},Y)$, i.e., ˜1.1, is akin to asking whether the puzzle is solvable from a given starting configuration. ˜1.2, on the other hand, asks how quickly the puzzle can be solved. In other words, given that the target configuration is reachable, how many steps are needed to reach it?
Connectivity in $\mathsf{FS}(\mathsf{Star}_{n},Y)$ has been extensively studied starting with the work of Wilson [13] and is well understood. However, prior to our work, no general bounds were known for the diameter of connected components in the graph. The main contribution of this work is to show that sliding puzzles with $n$ tiles, whenever solvable, can be solved in $\operatorname{poly}(n)$ steps. We also show how to generalize these results to other families of friends-and-strangers graphs.
Before we elaborate on our results, we describe previous work on friends-and-strangers graphs and related problems.
Connectivity in friends-and-stangers graphs.
Friends-and-strangers graphs were first introduced in their full generality and studied by Defant and Kravitz [6]. They studied connectivity for special cases such as $\mathsf{FS}(\mathsf{Path}_{n},Y)$ and $\mathsf{FS}(\mathsf{Cycle}_{n},Y)$, and derived necessary and sufficient conditions for $\mathsf{FS}(X,Y)$ to be connected. Subsequently, connectivity has been extensively studied for other families of friends-and-strangers graphs [2, 3, 5, 6, 9, 11, 12]. Stronger notions of connectivity, namely biconnectivity and $k$-connectivity, have also been studied [8].
Alon, Defant and Kravitz [2] initiated the study of extremal properties of friends-and-strangers graphs, including connectivity as a function of the minimum degrees of the graphs $X$ and $Y$ as well as when both $X$ and $Y$ are Erdős-Rényi random graphs. Bangachev [3] followed up with an almost complete characterization of connectivity as a function of the minimum degrees of the graphs $X$ and $Y$. Wang and Chen [12] and Milojević [9] further refined the connectivity results for random graphs.
The diameter of friends-and-strangers graphs.
The focus of our work is on ˜1.2, namely the diameter of friends-and-strangers graphs. Defant and Kravitz first proposed this as an open problem in their work [6, Section 7.3]. They noted that although in many settings $\mathsf{FS}(X,Y)$ is not connected, one may ask whether the diameter of each connected component in the graph is $\operatorname{poly}(n)$. (Observe that the size of $\mathsf{FS}(X,Y)$ is $n!$ – superexponential in $n$.) In other words, starting from any configuration, is it possible to get to any reachable configuration in a small number of steps?
Jeong [7] provided some partial answers to that question. He showed the existence of a family of friends-and-strangers graphs that contain connected components of diameter $e^{\Theta(n)}$. Jeong also investigated the diameter of some special classes of graphs; these results are reported in Table 1.
Token swapping.
Friends-and-strangers graphs are also a generalization of the token swapping problem that has been studied extensively in theoretical computer science. In this problem, $n$ distinct tokens are places on the vertices of an $n$-vertex graph $Y$. A move consists of swapping any two tokens whose locations are adjacent in $Y$. The graph of configurations corresponding to token swapping is then precisely given by $\mathsf{FS}(K_{n},Y)$, where $K_{n}$ is the complete graph on $n$ vertices. Yamanaka, Demaine, Ito, Kawahara, Kiyomi, Okamoto, Saitoh, Suzuki, Uchizawa, and Uno [14] first introduced the token swapping problem, and studied the complexity of finding the shortest path between any pair of configurations. They noted that if $Y$ is connected, then so is $\mathsf{FS}(K_{n},Y)$, and its diameter is always bounded by $O(n^{2})$. They accordingly asked whether the shortest path between any pair of configurations can be found in time polynomial in $n$. Miltzow, Narins, Okamoto, Rote, Thomas, and Uno [10] showed that this problem is NP-hard, and subsequent work has consequently studied the approximability of the problem [4].
1.1. Our Results
In this paper we provide new diameter bounds for a number of interesting families of friends-and-strangers graphs.
Our main technical result concerns $\mathsf{FS}(\mathsf{Star}_{n},Y)$, which models sliding puzzles. We show that any solvable sliding puzzle can be solved in $\operatorname{poly}(n)$ moves. In particular, the diameter of every connected component in $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is $O(n^{4})$. We also provide a family of $n$-vertex graphs $Y$ for which the diameter of $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is $\Omega(n^{3})$, showing that our upper bound is tight to within a factor of $O(n)$.
Theorem 1.3.
Let $Y$ be a graph on $n$ vertices. Then the diameter of any connected component of $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is $O(n^{4})$.
As mentioned previously, connectivity in $\mathsf{FS}(\mathsf{Star}_{n},Y)$ was fully characterized by Wilson [13]. However, Wilson’s approach is nonconstructive and based on properties of the symmetric and alternating groups. In contrast, our results are algorithmic and constructive, providing explicit short paths between pairs of configurations in $\mathsf{FS}(\mathsf{Star}_{n},Y)$. In fact, we make strong connections between the structure of $\mathsf{FS}(\mathsf{Star}_{n},Y)$ and $\mathsf{FS}(K_{n},Y)$, showing that paths in the latter can be converted at an additional multiplicative factor of $O(n^{2})$ to paths in the former. These results are summarized in Table 1.
Families Upper bound on Lower bound on
max possible diameter max possible diameter
General $e^{O(n\log n)}$ (trivial) $e^{\Omega(n)}$ [7]
$\mathsf{FS}(K_{n},Y)$ $2n^{2}-5n+3$ [7], $\binom{n}{2}$ Section˜3 $\binom{n}{2}$ Section˜3
$\mathsf{FS}(\mathsf{Cycle}_{n},Y)$ $8n^{4}(1+o(1))$ [7] $\binom{n}{2}$ [7]
$\mathsf{FS}(\mathsf{Path}_{n},Y)$ $\binom{n}{2}$ [7] $\binom{n}{2}$ [7]
$\mathsf{FS}(\mathsf{Star}_{n},\mathsf{Tree}_{n})$ $O(n)$ Section˜3 $\Omega(n)$ Section˜3
$\mathsf{FS}(\mathsf{Star}_{n},Y)$ $O(n^{4})$ Section˜4 $\Omega(n^{3})$ Section˜4
Table 1. Diameter bounds for friends-and-strangers graphs in prior work (black) and our work (blue). These bounds refer to the largest possible diameter for any connected component of the graph.
Our next result studies the diameter of $\mathsf{FS}(X,Y)$ when both $X$ and $Y$ are Erdős-Rényi random graphs. Our investigation is inspired by the work of Alon, Defant, and Kravitz [2], who established sharp thresholds for the probability $p$ to ensure the connectivity of $\mathsf{FS}(X,Y)$ when $X$ and $Y$ are both independently drawn from $\mathcal{G}(n,p)$. We focus on a single pair of permutations over the vertex set of $X$ and determine a condition on the edge probabilities $p$ and $q$ such that the pair is connected by a short path in $\mathsf{FS}(X,Y)$ when the graphs are drawn from $\mathcal{G}(n,p)$ and $\mathcal{G}(n,q)$ respectively. This result builds upon Theorem˜1.3.
Theorem 1.4.
Let $\tau$ and $\omega$ be arbitrary permutations over $[n]$. Let $X$ and $Y$ be random graphs over the vertex set $[n]$, independently drawn from $\mathcal{G}(n,p)$ and $\mathcal{G}(n,q)$ respectively, where $p$ and $q$ satisfy $pq\geq 100\log n/n$. Then, with probability at least $1-o(n^{-2})$, the distance between $\tau$ and $\omega$ in $\mathsf{FS}(X,Y)$ is $O(n^{6})$.
Alon, Defant and Kravitz studied the symmetric case where $p=q$. In order to prove connectivity of $\mathsf{FS}(X,Y)$, they required $p>p_{0}$ where $p_{0}^{2}=\operatorname{exp}(4(\log n)^{2/3})/n$. They conjectured that a threshold of $p^{2}=\Theta(n^{-1})$ should suffice for connectivity. Milojević [9] disproved this conjecture, showing that $p^{2}=\Omega(\log n/n)$ is necessary for connectivity. Wang and Chen [12] extended Alon et al.’s result to the asymmetric setting where the graphs $X$ and $Y$ have different edge probabilities $p$ and $q$. Their result also requires $pq\geq p_{0}^{2}$. In contrast, for our result a slightly smaller bound of $pq>O(1)\cdot\operatorname{exp}(\log\log n)/n$ suffices. This is because we bound the distance between a single pair of permutations with high probability. In particular, our result does not necessarily imply that $\mathsf{FS}(X,Y)$ is connected or has low diameter. We leave open the question of studying the diameter of $\mathsf{FS}(\mathcal{G}(n,p),\mathcal{G}(n,q))$.
Finally, we study diameter bounds for families of graphs $X,Y$ that have large minimum degree. This investigation is inspired by the work of Bangachev [3] who provided similar conditions for the connectivity of $\mathsf{FS}(X,Y)$.
Theorem 1.5.
Let $X$ and $Y$ be connected graphs on $n$ vertices such that
$\min(\delta(X),\delta(Y))+2\max(\delta(X),\delta(Y))\geq 2n.$
Then $\mathsf{FS}(X,Y)$ is connected and has diameter at most $O(n^{6})$.
Theorem 1.6.
Let $X$ and $Y$ be connected graphs on $n$ vertices such that $\delta(X)+\delta(Y)\geq\frac{3n}{2}$. Then $\mathsf{FS}(X,Y)$ is connected and has diameter at most 3n(n1)/23n(n-1)/2.
We can visualize these results, along with the results of Bangachev [3], in Figure˜1, where we identify pairs of values of $\delta(X)$ and $\delta(Y)$ for which $\mathsf{FS}(X,Y)$ is always connected, sometimes disconnected, or always has polynomial diameter. Observe that we prove a polynomial bound on the diameter for most of the regions where Bangachev [3] proved connectivity. The only exception is the green dotted region. This leads us to conjecture that whenever $\mathsf{FS}(X,Y)$ is connected, it has polynomially bounded diameter. Resolving this conjecture is the main problem left open by our work. See Section˜7 for a discussion of open questions.
Figure 1. Points $(\delta_{1}(n),\delta_{2}(n))$ for which $\mathsf{FS}(X,Y)$ must be connected (green), must be connected with polynomial diameter (blue, purple), or can be disconnected (red). Additive constants are omitted.
1.2. Organization of the paper
In Section˜2, we present some notation and basic claims which we will use throughout the rest of the paper. Then, in Section˜3, we bound the diameter of connected components of $\mathsf{FS}(X,Y)$ for special cases of $X$. In Section˜4, we prove our main theorem, Theorem˜1.3, and provide a lower bound on the maximum possible diameter of $\mathsf{FS}(\mathsf{Star}_{n},Y)$. In Section˜5, we prove Theorems˜1.5 and 1.6, and in Section˜6, we prove Theorem˜1.4. Finally, in Section˜7, we discuss some open directions for future work.
2. Preliminaries
2.1. Notation
In this subsection, we introduce the terminology which we use throughout this paper. For a graph $G$, we denote the vertex set of $G$ as $V(G)$ and the edge set of $G$ as $E(G)$.
In this paper, we will explore several families of graphs. Here are the most important ones with vertex set $n]=\{1,\dots,n\}$:
  • the complete graph, $K_{n}$, has an edge between any two vertices;
  • the path graph, $\mathsf{Path}_{n}$, has an edge between $i$ and $j$ when $j=i+1$;
  • the cycle graph, $\mathsf{Cycle}_{n}$, has an edge between $i$ and $j$ when $j=i+1\bmod{n}$;
  • the star graph, $\mathsf{Star}_{n}$, has an edge between $i$ and $j$ when $i$ or $j$ is equal to $1$.
We also define some other terms related to graph theory.
  • A cut vertex in a graph is a vertex such that when removed, the graph gains at least one connected component. A separable graph has at least one cut vertex. A graph is biconnected if it is connected but not separable.
  • A graph on $n$ vertices has a Hamiltonian cycle if it has a subgraph that is isomorphic to $\mathsf{Cycle}_{n}$.
  • A graph is $k$-regular if each vertex has degree $k$. A graph is regular if there exists some $k$ for which it is $k$-regular.
  • The diameter of a graph is the maximum distance between two vertices in a graph, where the distance between two vertices in a graph is defined as the length of the shortest path between the two vertices. A graph’s diameter is infinite if it is disconnected.
  • The minimum degree of a graph $G$, denoted as $\delta(G)$, is the infimum of the degrees of the vertices in $G$.
We now introduce notation related to friends-and-strangers graphs. By identifying $V(X)$ and $V(Y)$, we refer to a vertex of $\mathsf{FS}(X,Y)$ as a permutation mapping the vertices of $X$ onto the locations in $Y$. For a permutation $\sigma\in\mathsf{FS}(X,Y)$ and vertex $a\in V(X)$, we use $\sigma(a)\in V(Y)$ to denote the vertex in $Y$ that is located at $a$ in $\sigma$; Likewise, for a vertex $y\in V(Y)$, $\sigma^{-1}(y)$ denotes its location in $X$ under $\sigma$.
A move from a vertex to an adjacent vertex in $\mathsf{FS}(X,Y)$ is called an $(X,Y)$-friendly swap or, when $X$ and $Y$ are clear from the context, simply a friendly swap and mathematically denoted as a transposition. In particular, suppose $ab$ is an edge in $Y$ and $\sigma(a)\sigma(b)$ is an edge in $X$, then we represent the permutation resulting from the friendly swap of the elements $\sigma(a)$ and $\sigma(b)$ as $\sigma\circ(ab)$.
2.2. Preliminary observations
Here we collect some general results on graphs and permutations that we will use throughout the paper.
Lemma 2.1.
Any connected graph $G$ has a vertex which can be removed (along with its incident edges) such that the remaining graph is still connected.
Proof.
As $G$ is connected, it has some spanning tree $T$, which has some leaf $v$. Removing $v$, we see $T\setminus{v}$ is still a tree, in particular a spanning tree of $G\setminus{V}$, so $G\setminus{V}$ is still connected, as desired. ∎
Lemma 2.2.
Let $G$ be a biconnected graph and $e_{1}$ and $e_{2}$ be any two distinct edges in the graph. Then there is a simple (i.e., non-intersecting) cycle in $G$ containing $e_{1}$ and $e_{2}$.
Proof.
Let $e_{1}=(u_{1},v_{1})$ and $e_{2}=(u_{2},v_{2})$. Consider a graph $G^{\prime}$ obtained by replicating $G$; adding two new vertices $w_{1}$ and $w_{2}$; replacing the edge $(u_{1},v_{1})$ by the edges $(u_{1},w_{1})$ and $(w_{1},v_{1})$; and replacing the edge $(u_{2},v_{2})$ by the edges $(u_{2},w_{2})$ and $(w_{2},v_{2})$. Then, $G^{\prime}$ is also biconnected. Therefore, there exist two vertex disjoint paths from $w_{1}$ to $w_{2}$. These paths together form a cycle containing all four vertices $u_{1},u_{2},v_{1}$, and $v_{2}$. The same cycle in $G$ then contains the edges $e_{1}$ and $e_{2}$. ∎
3. Special Cases of Friends-and-Strangers Graphs
3.1. $\mathsf{FS}(K_{n},Y)$ for general $Y$
Recall from Table˜1 that Jeong [7] proved an upper bound of $2n^{2}-5n+3$ for the diameter of connected components of $\mathsf{FS}(K_{n},Y)$. We improve this bound to $\binom{n}{2}$ and show this is tight.
Theorem 3.1.
Let $Y$ be any graph on $n$ vertices. $\mathsf{FS}(K_{n},Y)$ is connected if and only if $Y$ is connected. Furthermore, every connected component of $\mathsf{FS}(K_{n},Y)$ has diameter less than or equal to $\binom{n}{2}$. Finally, $\mathsf{FS}(K_{n},\mathsf{Path}_{n})$ has a single connected component with diameter $\binom{n}{2}$.
Proof.
We will first assume that $Y$ is connected. We will show that $\mathsf{FS}(K_{n},Y)$ is connected and its diameter is bounded by $\binom{n}{2}$. Let $\sigma$ be our current permutation, and $\tau$ our target permutation. For any vertex $v$ in $Y$, we’ll define $f(v)=\sigma(\tau^{-1}(v))$. This is the current location of the vertex in $K_{n}$ that we want to move to $v$. Let $S\subset V(Y)$ be initialized as the empty set. Consider the following procedure, where we will maintain the invariant that the induced subgraph on $V(Y)\setminus S$ is connected:
  1. (1)
    Let $v$ be a vertex in $V(Y)\setminus S$ such that the induced subgraph on $V(Y)\setminus(S\cup\{v\})$ is connected, which always exists by Lemma˜2.1.
  2. (2)
    Then, there exists a path in the induced subgraph on $V(Y)\setminus S$ from $v$ to $f(v)$. Make $(K_{n},Y)$-friendly swaps until the person currently on $f(v)$ has moved to $v$.
  3. (3)
    Add $v$ to $S$, and repeat this procedure until $S=V(Y)$.
Note that at the end of this procedure, all people are in their intended positions, as once the person at $f(v)$ has moved to $v$, no friendly swaps are made involving $v$ for the rest of the procedure as henceforth $v\in S$. Finally, each iteration of our procedure takes a maximum of $n-\left\lvert S\right\rvert-1$ friendly swaps, meaning that the total number of friendly swaps made is at most $(n-1)+(n-2)+\cdots=\binom{n}{2}$.
If $Y$ is not connected, for a permutation $\sigma$ and a connected component $C$ in $Y$, let $\sigma^{-1}(C)$ denote the set of vertices of $K_{n}$ located on $C$. For two permutations $\sigma$ and $\tau$, we claim that $\sigma$ and $\tau$ are connected iff for all connected components $C$ in $Y$, $\sigma^{-1}(C)=\tau^{-1}(C)$. On the one hand, if the latter condition does not hold, then there exists a vertex $v$ in $K_{n}$ that is located at different components in $Y$ under $\sigma$ and $\tau$. But no friendly move can move the vertex from one component to another, so $\sigma$ and $\tau$ cannot be connected. On the other hand, if the condition holds, then we can transform $\sigma$ to $\tau$ by transforming the permutation over each connected component in sequence. The number of steps taken, when $Y$ has connected components $C_{1},\ldots,C_{k}$, is $\sum_{i\leq k}{{|C_{i}|}\choose 2}\leq{n\choose 2}$.
We will now show that equality holds in Theorem˜3.1 when $Y$ is isomorphic to $\mathsf{Path}_{n}$. We’ll number the vertices in $K_{n}$ $1,\cdots,n$, and number the vertices in $\mathsf{Path}_{n}$ $1,\cdots n$. Let $\sigma$ be the permutation $n,n-1,\cdots,1$ of the vertices of $K_{n}$ onto $\mathsf{Path}_{n}$, and let $\tau$ be the permutation $1,2,\ldots,n$. Define an inversion in a permutation $\omega$ to be a pair $(i,j)$ of vertices in $K_{n}$ such that $i>j$ but $\omega(i)<\omega(j)$.
Observe that the number of inversions in $\sigma$ is $\binom{n}{2}$. Each $(K_{n},\mathsf{Path}_{n})$-friendly swap decreases the number of inversions of our current permutation by at most $1$. Hence, if the people start in the configuration $\sigma$, it will take them at least $\binom{n}{2}$ friendly swaps to reach the configuration $\tau$, which has 0 inversions. ∎
3.2. $\mathsf{FS}(\mathsf{Star}_{n},Y)$ when $Y$ is a forest
The main result of this section is that when $Y$ is a forest, every connected component in $\mathsf{FS}(\mathsf{Star}_{n},Y)$ has linear diameter. We begin by proving this statement for trees.
Lemma 3.2.
Let $T$ be a tree on $n$ vertices. Then every connected component of $\mathsf{FS}(\mathsf{Star}_{n},T)$ is isomorphic to $T$.
Proof.
Let $s$ be the center vertex in $\mathsf{Star}_{n}$. Fix a starting configuration with $s$ on vertex $v$, and let the connected component of $\mathsf{FS}(\mathsf{Star}_{n},T)$ containing this permutation be $C$. We claim that there are precisely $n$ vertices in $C$, each of which has $s$ at a distinct vertex of $T$.
Suppose that we move $s$ to another vertex $w$ through some sequence of friendly swaps. Observe that if at any point we move $s$ from some vertex $a$ to $b$ and then back from $b$ to $a$, the configuration has not changed (we have essentially done and undone an action). Hence, if $s$ is moved to $w$ in a non-simple path, then we can delete some friendly swaps from this sequence of moves. Hence, we may assume that $s$ is moved to $w$ in a simple path. Because there is exactly one simple path from $v$ to $w$ in $T$, we conclude that there is exactly one configuration with $s$ at $w$ in $C$. This shows that the number of vertices in $C$ is $n$.
To finish the claim, observe that two configurations in $C$ are adjacent if and only if the positions of $s$ in these configurations are adjacent. Hence, it follows that $C$ is isomorphic to $T$, with a natural bijection $T\to C$ mapping any vertex $v$ in $T$ to the position in $C$ with $s$ at $v$. ∎
We obtain the bound for forests as a simple corollary.
Theorem 3.3.
Let $F$ be a forest on $n$ vertices. Then, the connected components of $\mathsf{FS}(\mathsf{Star}_{n},F)$ have diameter at most n1n-1.
Proof.
Again letting $s$ denote the central vertex in $\mathsf{Star}_{n}$, consider an arbitrary starting position $\sigma\in V(\mathsf{FS}(\mathsf{Star}_{n},F))$, viewed as a permutation of people from $\mathsf{Star}_{n}$ onto positions from $F$. Observe that if $C$ is a connected component of $F$ which $s$ is not in, then the people on $C$ cannot be moved. Hence, if $C^{\prime}$ is the connected component which $s$ is placed in, then by Lemma˜3.2, the connected component containing $\sigma$ must be isomorphic to $C^{\prime}$, so has diameter at most $\left\lvert C^{\prime}\right\rvert-1\leq n-1$. ∎
$3.$ $\mathsf{FS}(\mathsf{Star}_{n},Y)$ when $Y$ is a cycle
In this section, we’ll show that the diameters of the connected components in $\mathsf{FS}(\mathsf{Star}_{n},\mathsf{Cycle}_{n})$ are at most $\binom{n}{2}$, and that, more strongly, each connected component is isomorphic to $\mathsf{Cycle}_{n(n-1)}$. Furthermore, we’ll fully characterize the configurations in each connected component of $\mathsf{FS}(\mathsf{Star}_{n},\mathsf{Cycle}_{n})$, a result that is valuable in Section˜4. The first two results can be found in [7], but for completeness, we prove them below. Let $\sigma$ denote the current bijection from the vertices in $\mathsf{Star}_{n}$ onto the vertices in $\mathsf{Cycle}_{n}$.
Theorem 3.4.
Each connected component of $\mathsf{FS}(\mathsf{Star}_{n},\mathsf{Cycle}_{n})$ is isomorphic to $\mathsf{Cycle}_{n(n-1)}$. More strongly, if $s$ is the center of $\mathsf{Star}_{n}$, $a$ is an arbitrary vertex in $\mathsf{Star}_{n}$, and $x$ and $y$ are arbitrary vertices in $\mathsf{Cycle}_{n}$, then there is exactly one configuration in any connected component of $\mathsf{FS}(\mathsf{Star}_{n},\mathsf{Cycle}_{n})$ for which $\sigma(s)=x$ and $\sigma(a)=y$.
Proof.
We’ll prove the second fact, after which the first fact easily follows. Fix an orientation of the vertices of $\mathsf{Cycle}_{n}$ (which we’ll call “clockwise”). Observe that the clockwise ordering of the vertices of $\mathsf{Star}_{n}$ except for $s$ around $\mathsf{Cycle}_{n}$ cannot change, because none of these vertices can cross each other. In other words, given the positions of $s$ and $a$ on $\mathsf{Cycle}_{n}$, the positions of every other vertex in $\mathsf{Star}_{n}$ is fixed.
We now show that $s$ and $a$ can be moved to any two positions through a series of friendly swaps. As in the theorem statement, let $x$ be the target position of $s$, and $y$ be the target position of $a$ in $\mathsf{Cycle}_{n}$. Observe that if we friendly swap $s$ in a clockwise direction with a vertex other than $a$, then $a$ will not move. On the other hand, if we friendly swap $s$ in a clockwise direction with $a$, then $a$ moves counterclockwise by one vertex. Hence, if we repeatedly friendly swap $s$ in a counterclockwise direction, $a$ will eventually move to $y$. Then, we can friendly swap $s$ in the path formed by the induced subgraph of $V(G)\setminus y$ until it is on $x$.
We have shown that our connected component has n(n1)n(n-1) vertices. Furthermore, because each vertex has degree $2$ (because $s$ can only be friendly swapped in two directions), it follows that our connected component is isomorphic to $\mathsf{Cycle}_{n(n-1)}$. ∎
In later sections, we will use the fact that by moving $s$ around a cycle $C$ in a graph $Y$, we can move any set of $k$ consecutive vertices on $C$ to any set of $k$ consecutive positions in $C$. This follows easily by the previous theorem.
4. $\mathsf{FS}(\mathsf{Star}_{n},Y)$ for general $Y$
In this section we will prove our main theorem for star graphs, which we restate below for convenience.
See 1.3
4.1. Overview of proof
Throughout this section we will use $s$ to denote the vertex at the center of $\mathsf{Star}_{n}$. As a default, we’ll assume that configurations in $\mathsf{FS}(\mathsf{Star}_{n},Y)$ are permutations from $V(\mathsf{Star}_{n})$ to $V(Y)$. Note that any friendly swap in $\mathsf{FS}(\mathsf{Star}_{n},Y)$ from a starting permutation $\sigma$ implements a transposition $(sx)$ for some $x\in V(\mathsf{Star}_{n})$ such that $\sigma(s)$ and $\sigma(x)$ are adjacent in $Y$. We will also consider sequences of moves along paths in $Y$. For any permutation $\sigma$ and any simple path $P$ in $Y$ starting at $\sigma(s)$ (or simple cycle containing $\sigma(s)$) and containing vertices $\sigma(p_{1}),\sigma(p_{2}),\ldots,\sigma(p_{k})$ in that order, we can implement the rotation $(sp_{1}p_{2}\ldots p_{k})$ through a sequence of $k<n$ friendly swaps. We will use $\left\lvert P\right\rvert$ to denote the number of edges $k$ in such a path $P$. Finally, for a path $P$ in $Y$, we use $\operatorname{rev}(P)$ to denote its reversal – the sequence of edges in $P$ in backwards order.
Our proof is based on identifying subgraphs of the graph $Y$ that are isomorphic to a $\theta$-graph, defined below, and establishing moves between permutations over those subgraphs. For ease of exposition, we use a slightly non-standard notation for $\theta$-graphs that we will now introduce.
Definition 4.1.
$\theta(i,j,k)$ is a graph over $i+j+k+2$ vertices, of which two vertices that we call the endpoints have degree $3$ and the remaining have degree $2$. The endpoints, say $v$ and $w$, are connected by three vertex-disjoint paths: $v,x_{1},x_{2},\ldots,x_{i},w$, $v,y_{1},y_{2},\ldots,y_{j},w$, $v,z_{1},z_{2},\ldots,z_{k},w$. Without loss of generality we assume $i\leq j\leq k$. Figure 2 illustrates $\theta(1,2,2)$.
Figure 2. The graph $\theta(1,2,2)$.
Fact 4.2.
Every biconnected graph that is not a cycle contains a $\theta$-subgraph, i.e., a subgraph isomorphic to a $\theta$-graph.
We use the following connectivity result from Wilson [13].
Theorem 4.3 ([13]).
Let $Y$ be a biconnected graph on $n$ vertices which is not isomorphic to $\mathsf{Cycle}_{n}$ or $\theta(1,2,2)$. Then, if $Y$ is bipartite, $\mathsf{FS}(\mathsf{Star}_{n},Y)$ has exactly $2$ connected components. Vertices are placed in these connected components based on the parity of the permutation without $s$. Otherwise, $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is connected.
The bulk of our proof focuses on graphs $Y$ that are biconnected. We then combine this case with our analysis of forests to obtain the main result. The analysis of biconnected graphs identifies two elementary sequences of moves in $\mathsf{FS}(\mathsf{Star}_{n},Y)$, which can be concatenated together to obtain short paths between any pair of connected permutations. The following lemmas formalize these moves.
Lemma 4.4.
[…]
Lemma 4.5.
[…]
Lemma 4.6.
[…]
The rest of this section is organized as follows. We will first prove Theorem 1.3 using Lemmas 4.4 and 4.5. Then in Section 4.3 we will prove Lemmas 4.4, 4.5, and 4.6 using $\theta$-subgraphs as building blocks. Finally, in Section 4.4 we will use a case analysis to define and prove the existence of the 3-cycles in permutations on $\theta$-graphs that support our entire argument.
4.2. Proof of Theorem 1.3
We first prove the special case of the theorem for biconnected graphs, using Lemmas 4.4, 4.5, and 4.6.
Lemma 4.7.
Let $Y$ be a biconnected graph on $n$ vertices. Then the diameter of any connected component of $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is $O(n^{4})$.
Proof.
Note that when $Y=\theta(1,2,2)$, $\mathsf{FS}(\mathsf{Star}_{7},Y)$ has six connected components, each with $O(1)$ diameter. Furthermore, we show in Theorem˜3.4 that when $Y=\mathsf{Cycle}_{n}$, the diameter of each component of $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is $O(n^{2})$. In the rest of the proof we exclude these cases.
Consider any two permutations $\sigma,\omega$ that are in the same connected component in $\mathsf{FS}(\mathsf{Star}_{n},Y)$. First, we observe that we may assume $\sigma(s)=\omega(s)$ without loss of generality. If not, we can find a simple path $P$ in $Y$ from $\sigma(s)$ to $\omega(s)$ and move $s$ to $\omega(s)$ along $P$ in $\left\lvert P\right\rvert<n$ steps.
We now consider two cases. First suppose $Y$ is bipartite. Assume without loss of generality that $\omega=12\cdots n$ by relabeling $V(Y)$ as necessary so that the number of inversions in $\omega$ is 0. We claim that $\sigma$ and $\omega$ have the same parity. This is because in any path from $\sigma$ to $\omega$ in $\mathsf{FS}(\mathsf{Star}_{n},Y)$, each friendly swap moves the location of $s$ in $Y$, $s$ starts and ends in the same location, and $Y$ is bipartite; thus, the number of swaps in any path from $\sigma$ to $\omega$ is even, and applying an arbitrary transposition to a permutation changes its number of inversions by an odd number.
Next, define $Y^{\prime}$ to be the induced subgraph of the vertex set $V(Y)\setminus\{\sigma(s)\}$. Because $Y$ is biconnected, it follows that $Y^{\prime}$ is connected. Now, let $S$ be any sequence of friendly swaps of length $O(n^{2})$ in $\mathsf{FS}(K_{n-1},Y^{\prime})$ which starts at $\sigma$ and ends at $\omega$, which exists because $Y^{\prime}$ is connected. Then, $\left\lvert S\right\rvert$ must be even because $\sigma$ and $\omega$ have the same parity. We can therefore pair up the first and second moves in $S$, the third and fourth, and so on. We’ll denote each pair of friendly swaps as good if the two edges in $Y$ that we have swapped over share a vertex, and bad otherwise.
We first claim that there exists a path $S$ from $\sigma$ to $\omega$ in $\mathsf{FS}(K_{n},Y)$ that contains $O(n^{2})$ good pairs and $O(n)$ bad pairs. This easily follows by the construction in our proof of Theorem˜3.1, which provides a path $S$ of length $O(n^{2})$. In particular, observe that our algorithm consists of consecutively making many moves which swap one fixed vertex into its desired position. We do this $O(n)$ times. Observe that a bad pair can only occur in between these consecutive sequences, when one sequence finishes in the first swap and another sequence starts in the second swap. Hence, there can clearly only be $O(n)$ bad pairs.
To finish, we claim that we can perform each pair of swaps in $S$ in $O(n^{2})$ friendly swaps if it’s good, and $O(n^{3})$ if it is not. The latter case follows easily from Lemma˜4.6. For the former case, observe that performing two friendly swaps over edges that share a vertex is equivalent to performing a move of the form $(abc)$ for some vertices $a,b,c$, because all three vertices in the edges do not return to their starting positions. Hence, we can invoke Lemma˜4.5 to solve this case.
To finish, concatenating these sequences of friendly swaps yields us a sequence of a total of $O(n^{3}\cdot n+n^{2}\cdot n^{2})=O(n^{4})$ friendly swaps. This finishes the proof of this case.
Finally, suppose that $Y$ is not bipartite and not isomorphic to $\theta(1,2,2)$ or $\mathsf{Cycle}_{n}$. In this case, we will show that if $\sigma$ and $\omega$ are in the same connected component in $\mathsf{FS}(K_{n},Y)$, then they are at distance in $O(n^{4})$ in $\mathsf{FS}(\mathsf{Star}_{n},Y)$. Since $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is a subgraph of $\mathsf{FS}(K_{n},Y)$, this implies the lemma. Similarly to the previous case, let $S$ be a sequence of moves in $\mathsf{FS}(K_{n},Y)$ which moves from $\sigma$ to $\omega$. We claim that we can perform any move in $S$, $(ab)$, in $O(n^{2})$ friendly swaps. Because the length of $S$ is $O(n^{2})$, this will finish.
Let the current permutation be $\sigma$, and consider a friendly swap $(ab)$ in $\mathsf{FS}(K_{n},Y)$ that takes us to $\sigma\circ(ab)$. Let $x$ and $y$ be the adjacent vertices in $Y$ that satisfy Lemma˜4.4. Since $(ab)$ is a friendly swap in $\mathsf{FS}(K_{n},Y)$, $\sigma(a)\sigma(b)$ is also an edge in $Y$. By Lemma˜2.2, there exists a cycle $C$ through $\sigma(a),\sigma(b),x$, and $y$. We conclude the proof by splitting into three separate cases.
  1. (1)
    First, suppose that $\sigma(s)$ is on this cycle $C$. Then we move $s$ around this cycle until $a$ and $b$ are on $e$ and $f$; a total of $O(n^{2})$ moves. Then, we apply the transposition $(ab)$ as guaranteed by Lemma˜4.4, switching the positions of $a$ and $b$. Finally, we revert the moves of $s$ along the cycle, which restores our initial position, except with the positions of $a$ and $b$ switched.
  2. (2)
    On the other hand, if $s$ is not on the cycle $C$, consider any simple path in $Y$ from $\sigma(s)$ to $C$ that does not contain $\sigma(a)$ or $\sigma(b)$. If such a path exists, we move $s$ to $C$ along this path, implement the $O(n^{2})$ moves described earlier, and then revert the path restoring the initial position with the positions of $a$ and $b$ switched.
  3. (3)
[…]
Each of the cases involves a sequence of $O(n^{2})$ moves to reach $\sigma\circ(ab)$, completing the proof. ∎
We now extend our argument to non-biconnected graphs $Y$. We do this by inductively breaking up the graph along its cut vertices. The following lemma establishes the inductive step.
Lemma 4.8.
Let $G$ be a connected graph with a cut vertex $v$. Let $X$ and $Y$ be the subgraphs created when $G$ is split at $v$, each containing $v$. Then,
$\operatorname{diam}(\mathsf{FS}(G,\mathsf{Star}_{|V(G)|}))\leq\max(\operatorname{diam}(X),\operatorname{diam}(Y))$ $+\operatorname{diam}(\mathsf{FS}(X,\mathsf{Star}_{|V(X)|}))$
$+\operatorname{diam}(\mathsf{FS}(Y,\mathsf{Star}_{|V(Y)|})).$
Proof.
[…]
  1. (1)
    Suppose that $s$ is in the same subgraph (without loss of generality, $X$) in both $\sigma$ and $\tau$. In this case, we perform the following operation: Move $s$ to $v$ in at most $\operatorname{diam}(X)$ friendly swaps. Then, perform at most $\operatorname{diam}(\mathsf{FS}(Y,\mathsf{Star}_{|V(Y)|}))$ to put every person on $Y$ in their position in $\tau$ (note that $s$’s intended position is defined as $v$ here). Finally, perform at most $\operatorname{diam}(\mathsf{FS}(X,\mathsf{Star}_{|V(X)|}))$ friendly swaps to put every person on $X$ in their position in $\tau$. This algorithm clearly satisfies the bound.
  2. (2)
    If $s$ is in $X$ in $\sigma$ but in $Y$ in $\tau$, perform the following operations: perform at most $\operatorname{diam}(\mathsf{FS}(X,\mathsf{Star}_{|V(X)|}))$ friendly swaps to put every person on $X$ in their position in $\tau$ (again with $s$’s intended position defined to be $v$ here), and perform $\operatorname{diam}(\mathsf{FS}(Y,\mathsf{Star}_{|V(Y)|}))$ friendly swaps to put every person on $Y$ in their position in $\tau$. This algorithm clearly satisfies the bound. ∎
We are now ready to prove our main theorem using Lemmas˜4.7 and 4.8.
Proof of Theorem˜1.3.
We prove this by strong induction on $n$. Our base cases are $n=1$ and $n=2$, for which the diameter of any connected component of $\mathsf{FS}(G,\mathsf{Star}_{n})$ is clearly at most $2!=2$. We now continue with our inductive step. Assume that for all $k<n$, the bound $\operatorname{diam}(\mathsf{FS}(G,\mathsf{Star}_{k}))\leq dk^{4}$ holds for some constant $d$, and that for all $k$, the bound $\operatorname{diam}(\mathsf{FS}(G,\mathsf{Star}_{k}))\leq dk^{4}$ holds for biconnected $G$. Note that a value of $d$ satisfying this must exist by Lemma˜4.7, and we may assume $d\geq 1$. We will show that this bound holds for $k=n$. We split into the following three cases.
  1. (1)
    If $G$ is biconnected, then we are done by Lemma˜4.7.
  2. (2)
    If $G$ is disconnected and $s$ is in connected component $C$, then the diameter of the current connected component of $\mathsf{FS}(G,\mathsf{Star}_{n})$ is equal to the diameter of a connected component of $\mathsf{FS}(C,\mathsf{Star}_{|C|})$, which is at most $d\left\lvert C\right\rvert^{4}\leq dn^{4}$.
  3. (3)
    If $G$ is connected but has a cut vertex $v$, then we can use Lemma˜4.8. Let the subgraphs created by splitting $G$ at $v$ be $X$ and $Y$, with $p$ and np+1n-p+1 vertices, respectively. Clearly, each of $X$ and $Y$ must have at least $2$ vertices, so by Lemma˜4.8, the maximum possible value of our diameter is thus $d(2^{4}+(n-1)^{4})+n-1\leq d(2^{4}+(n-1)^{4}+n-1)$. The term $2^{4}+(n-1)^{4}+n-1$ can be easily checked to be less than $n$ for $n\geq 3$. Hence, the diameter of $\mathsf{FS}(\mathsf{Star}_{n},G)$ is less than or equal to $dn^{4}$. This completes our inductive step. ∎
4.3. Analysis of elementary moves in general biconnected graphs
We will now prove Lemmas 4.4, 4.5, and 4.6. These proofs utilize the existence of moves of the form $(xyz)$ in some $\theta$-subgraph of the graph $Y$. These moves are critical to understanding the structure of $\mathsf{FS}(\mathsf{Star}_{n},Y)$. The following lemma, which we will prove in Section˜4.4, shows the existence of these moves.
Lemma 4.9.
Let $H=\theta(i,j,k)$ be a $\theta$-graph on $n=i+j+k+2$ vertices such that $(i,j,k)\neq(1,2,2)$. Then there exist vertices $x,y,z$ in $H$ such that
  • xyxy and $yz$ are edges in $H$,
  • $y$ has degree $2$ in $H$, and,
[…]
We first need the following lemma to help prove Lemma˜4.4.
Lemma 4.10.
Let $Y$ be a biconnected nonbipartite graph such that $Y$ contains a $\theta$-subgraph and any $\theta$-subgraph of $Y$ that contains an odd cycle is isomorphic to $\theta(1,2,2)$. Then $Y$ is isomorphic to $\theta(1,2,2)$.
Proof.
Assume the contrary, that $Y$ only contains $\theta$-subgraphs isomorphic to $\theta(1,2,2)$ but is not isomorphic to $\theta(1,2,2)$ itself. Let $H$ be a $\theta$-subgraph of $Y$ that is $\theta(1,2,2)$. We will show in our proof that $Y$ contains a $\theta$-subgraph which is not isomorphic to $H$. Because all pairs of two cycles in $H$ share at least two edges, in many cases it suffices to show that there exists a $\theta$-subgraph of $Y$ with two cycles that only share one edge.
Let $C$ be a cycle in $Y$ that is not completely contained in $H$ but shares an edge with $H$. Let $P$ be the maximal path in $C$ such that none of $P$’s edges are in $H$. Let $x$ and $y$ be the endpoints of $P$, which are in $H$.
Observe that any graph isomorphic to $H$ has only $5$-cycles and $6$-cycles. Let the $6$-cycle in $H$ be $D$, and let the two $5$-cycles be $E$ and $F$. Let the maximal path shared by $D$ and $E$ be $P_{1}$ and the maximal path shared by $D$ and $F$ be $P_{2}$. We first claim that $x$ and $y$ must be on $D$. If not, then one of $x$ and $y$ is in the center vertex of our $\theta$-graph, and the minimal path $Q$ between $x$ and $y$ can only go through the edges of at most one of $P_{1}$ and $P_{2}$. Assume without loss of generality that the path does not pass through any edges of $P_{2}$. Then, the cycle formed by the edges of $P$ and $Q$ shares only one edge with $F$, and therefore the $\theta$-graph formed by the union of the edges of this cycle and $F$ cannot be isomorphic to $H$, contradiction.
We next claim that $x$ and $y$ must be of distance $3$ in $D$. This is because if this were not true and $Q$ and $R$ were the paths in $D$ when it is split by $x$ and $y$, then $\left\lvert\left\lvert Q\right\rvert-\left\lvert R\right\rvert\right\rvert\geq 2$ because both paths are not the same length and their lengths sum up to $6$. Thus, the number of vertices in the cycles $C_{1}=P\cup Q$ and $C_{2}=P\cup R$ differ by at least $2$ (this difference is the same as the previous mentioned value), meaning that one of these two cycles does not have $5$ or $6$ vertices. Assume without loss of generality that this cycle is $C_{1}$. Observe that $C_{1}$ must share an edge with $D$, and thus one of $E$ and $F$. Assume without loss of generality that this cycle is $E$. Then, the $\theta$-graph formed by the union of $C_{1}$ and $E$ contains an odd cycle ($E$) and a cycle of size not equal to $5$ or $6$, so it cannot be isomorphic to $H$. The argument is exactly the same when $C_{1}$ shares a cycle with $F$.
Now that we know $x$ and $y$ must be of distance $3$ in $D$, we proceed with the rest of our proof. Let $a$ and $b$ be the endpoints of $H$. Then, we assume without loss of generality that $x$ is adjacent to $a$ and on $E$ and $y$ is adjacent to $b$ and on $F$. Let the path from $x$ to $y$ that is on $D$ and passes through $a$ be $R$, and let $C^{\prime}$ be the cycle formed by the paths $P$ and $R$. Clearly $C^{\prime}$ and $E$ share exactly one edge, namely $ae$. This implies that the $\theta$-graph formed by the union of $C^{\prime}$ and $E$, which contains an odd cycle, cannot be isomorphic to $\theta(1,2,2)$. This is our final contradiction. ∎
We now prove Lemma˜4.4, which we restate for convenience. See 4.4
Proof.
Since $Y$ isn’t isomorphic to a cycle, it follows that $Y$ must have at least two cycles, and because $Y$ is biconnected, it must have two cycles which share an edge. Hence, $Y$ contains a $\theta$-subgraph. Because $Y$ is not isomorphic to $\theta(1,2,2)$, it follows from Lemma 4.10 that $Y$ contains a $\theta(i,j,k)$ subgraph $H$ with an odd cycle that is not isomorphic to $\theta(1,2,2)$. Because $H$ has an odd cycle, it follows that one of $i,j,k$ is odd, but not all three. Therefore, $H$ contains two odd cycles.
Lemma 4.9 provides us with vertices $x,y,z$ in $H$ such that for all permutations $\sigma$ with $\sigma(s)\in V(H)$, we have $\sigma|_{H}$ and $\sigma|_{H}\circ\sigma|_{H}(xyz)$ at distance $O(n)$ in $\mathsf{FS}(\mathsf{Star}_{n},Y)$. Furthermore, two out of the three cycles in $H$ contain these three vertices. Because two out of the three cycles in $H$ also have an odd number of edges, this means that one of the cycles, call it $C$, is odd and contains $x$, $y$ and $z$.
[…]
Let the current circular permutation of the vertices of $\mathsf{Star}_{n}$ (except for $s$) on $C$ be $c_{1}c_{2}\ldots c_{k}$ such that $c_{1}=\hat{x}$ and $c_{2}=\hat{y}$. Observe that $k$ is even. Perform the move $(c_{1}c_{2}c_{3})$ so that $c_{1}$ moves to $z$, $c_{2}$ moves to $x$, and $c_{3}$ moves to $y$. We can view this as moving $c_{1}$ back two spaces in our circular permutation. Then, move $s$ twice around $C$ so that $c_{1}$ ends up back on $x$. We then repeat this process a total of $\frac{k}{2}$ times. Because $c_{1}$ is moved back two places in the circular permutation in each iteration, it has been moved back a total of $k$ places. Because there are k1k-1 elements in our circular permutation other than $c_{1}$, it follows that $c_{1}$ ends up one space behind its original spot. Hence, our new circular permutation is $c_{2}c_{1}\ldots c_{k}$. We now move $s$ back around the cycle until $c_{2}$ is on $x$ and $c_{1}$ is on $y$. The total number of moves this takes is $O(n^{2})$, because each move of the form $(\sigma^{-1}(x)\sigma^{-1}(y)\sigma^{-1}(z))$ takes $O(n)$ friendly swaps.
We now need to handle the case where $\sigma(s)\not\in C$. In this case, we consider a simple path from $\sigma(s)$ to $C$ in $Y$ that does not contain $x$ or $y$. If such a path exists, we move $s$ to $C$ along this path, follow the aforementioned sequence of moves to swap the elements $\hat{x}$ and $\hat{y}$ and revert the sequence of moves along the path to restore our initial position with $\hat{x}$ and $\hat{y}$ swapped.
Finally, if the only paths from $\sigma(s)$ to $C$ end in $x$ and $y$, we perform a longer sequence of moves to move $s$ to $C$ while maintaining the positions of $\hat{x}$ and $\hat{y}$ at $x$ and $y$ respectively; then swap $\hat{x}$ and $\hat{y}$ as described previously; and then restore our initial position by reversing the sequence of moves.
The sequence of moves is described as follows. Let $C^{\prime}$ be the cycle formed by the paths from $\sigma(s)$ to $x$ and $y$ along with the edge $xy$. We first move $s$ around this cycle to move $\hat{x}$ and $\hat{y}$ to adjacent vertices on the path $C^{\prime}\setminus\{xy\}$. We then move $s$ around $C\cup C^{\prime}\setminus\{xy\}$ to move $\hat{x}$ and $\hat{y}$ to adjacent vertices on the path $C\setminus\{xy\}$ with $s$ ending up at a vertex in $C$. Finally, we move $s$ along the cycle $C$ until $\hat{x}$ and $\hat{y}$ move back to $x$ and $y$. This completes the procedure of moves. It can be seen easily that each subsequence involves at most $O(n^{2})$ moves. ∎
We’ll now prove Lemmas˜4.5 and 4.6. First, we need the following auxiliary lemma.
Lemma 4.11.
Let $Y$ be a biconnected graph on $n$ vertices, and let $\sigma:V(\mathsf{Star}_{n})\to V(Y)$ be a permutation in $\mathsf{FS}(\mathsf{Star}_{n},Y)$. Let $x$, $y$, and $z$ be distinct vertices in $\mathsf{Star}_{n}$, where $s\not\in\left\{x,y,z\right\}$, and let $u$, $v$, and $w$ be distinct vertices in $Y$ such that $\sigma(y)$ is adjacent to $\sigma(x)$ and $\sigma(z)$ in $Y$, and $v$ is adjacent to $u$ and $w$. Then, there exists a series of friendly swaps of length $O(n^{2})$ in $\mathsf{FS}(\mathsf{Star}_{n},Y)$ which moves the set $\{x,y,z\}$ to $\{u,v,w\}$.
Before proving this lemma, we first show how it proves Lemmas˜4.5 and 4.6, which we restate for convenience.
See 4.5
Proof.
Let $H$ be a $\theta$-subgraph of $Y$, which exits by ˜4.2. Let $x,y,z$ be the vertices described in Lemma 4.9 for $H$.
[…]
  1. (1)
    Using Lemma 4.11, we first claim that through a sequence $S$ of $O(n^{2})$ friendly swaps, we can move the elements $\{\hat{a},\hat{b},\hat{c}\}$ to $\{x,y,z\}$ in some order.
  2. (2)
[…]
  3. (3)
[…]
The entire procedure takes $O(n^{2})$ friendly swaps, so we are done. ∎
See 4.6
Proof.
We obtain this result by applying Lemma˜4.5 repeatedly. Observe that by Lemma˜2.2, there exists a cycle through the edges $ab$ and $cd$.
Because of this, there exists a path that starts with edge $ab$ and ends with edge $cd$. We denote this path as $P$ and assume via renaming vertices that the endpoints of $P$ are $a$ and $d$.
Suppose that the current permutation of the vertices in $\mathsf{Star}_{n}$ on $P$ is $\hat{a},\hat{b},x_{1},\dots,x_{k},\hat{c},\hat{d}$. We apply Lemma˜4.5 to implement the move $(\hat{a}\hat{b}x_{1})$ in $O(n^{2})$ steps. This results in the permutation $\hat{b},x_{1},\hat{a},x_{2},\ldots,x_{k},\hat{c},\hat{d}$ of the vertices on $P$. We then apply Lemma˜4.5 to implement the move $(x_{1}\hat{a}x_{2})$ in $O(n^{2})$ steps. This results in the permutation $\hat{b},\hat{a},x_{2},x_{1},\ldots,x_{k},\hat{c},\hat{d}$ of the vertices on $P$. We proceed in this manner a total of $k+2\leq n$ times. The final permutation obtained is $\hat{b},\hat{a},x_{1},x_{2}\ldots,x_{k},\hat{d},\hat{c}$, and we are done. ∎
To conclude this subsection we prove Lemma˜4.11.
Proof of Lemma˜4.11.
Our first claim is that there exists $k<4$ and a sequence of cycles $C_{1},C_{2},\ldots,C_{k}$ such that $C_{1}$ contains the edges $\sigma(x)\sigma(y)$ and $\sigma(y)\sigma(z)$ while $C_{k}$ contains the edges $uv$ and $vw$ and, for each $i<k$, $C_{i}$ and $C_{i+1}$ share at least one edge.
To construct this, we first choose cycles $C_{1}$ and $C_{k}$ that satisfy our requirements, which exist by Lemma˜2.2. Then, if $C_{1}$ and $C_{k}$ do not share an edge (otherwise we set $k=2$ and are done), we can choose two arbitrary edges $e_{1}$ in $C_{1}$ and $e_{2}$ in $C_{k}$, and find a cycle $C_{2}$ containing these edges (again using Lemma˜2.2). This yields the sequence $C_{1},C_{2},C_{k}$ and $k=3$.
We now are ready to describe a sequence of friendly swaps which result in $\{x,y,z\}$ being mapped to $\{u,v,w\}$. If $k\geq 2$, we will first describe a procedure to move $x,y,$ and $z$ to $C_{2}$ while maintaining their adjacency, i.e., $\left\{x,y,z\right\}$ forms a path in $C_{2}$, though not necessarily in that order. Then, this algorithm clearly also applies to moving $x,y,$ and $z$ from $C_{2}$ to $C_{3}$, while maintaining their adjacency, if $k=3$. We have two key steps.
First, we show how to move $s$ onto $C_{1}$ while keeping $\left\{x,y,z\right\}$ a path on $C_{1}$. If $s$ is already on $C_{1}$, then we are done; otherwise, because $Y$ is biconnected, it follows that there are two disjoint paths from $s$ to $C_{1}$ that end at different vertices. This is because if these two paths didn’t exist, then there exists a vertex $\rho$ such that every path from $s$ to $C_{1}$ ends at $\rho$, where we assume that our paths stop immediately upon reaching $C_{1}$. This implies that when $\rho$ is removed, $Y$ is disconnected, contradicting $Y$ being biconnected. From now on, we’ll assume that our two paths from $s$ to $C_{1}$ stop right when they reach $C_{1}$. If either of these two vertices is not $\sigma(x),\sigma(y),$ or $\sigma(z)$ (which we will denote as $q,r,$ and $t$ for simplicity), then we are done; we simply move $s$ along the corresponding path. Otherwise, we have a few cases.
  1. (1)
    Our first case is when $C_{1}$ has more than three vertices and the two paths $P_{1}$ and $P_{2}$ end at $q$ and $r$, respectively. Then we swap $s$ along $P_{2}$ to $r$ and then to $q$ so that $x$ is now on $r$ and $y$ is on the vertex of $P_{2}$ adjacent to $r$. Then, observe that the cycle $C^{\prime}$ formed by $(E(C_{1})\cup E(P_{1})\cup E(P_{2}))\setminus\left\{qr\right\}$ contains $x,y,z$, and $s$, and $yx$ and $xz$ are edges. Finally, we can move $s$ around $C^{\prime}$ until $x,y,z,$ and $s$ are all in $C_{1}$. Then we’re done.
    Note that an essentially identical argument also addresses the case when $C_{1}$ has more than three vertices and the two paths $P_{1}$ and $P_{2}$ end at $r$ and $t$.
  2. (2)
    Our second case is when $C_{1}$ has more than four vertices and two paths $P_{1}$ and $P_{2}$ end at $q$ and $t$, respectively. We follow a similar procedure. We first move $s$ through $P_{2}$ to $t$ and back through $r$, $q$, and $P_{1}$, followed by moving $s$ again through $P_{2}$ to $t$ and then $r$ and $q$. After this, $x$ is on $t$, and $y$ and $z$ are in $P_{1}\cup P_{2}$. Furthermore, $x$ and $y$ are still in adjacent positions, and so are $y$ and $z$. Hence, the cycle $C^{\prime}$ formed by $(E(C_{1})\cup E(P_{1})\cup E(P_{2}))\setminus\{qr,rt\}$ contains $x,y,z$, and $s$, and $xy$ and $yz$ are edges. Finally, we can move $s$ around $C^{\prime}$ until $x,y,z,s$ are all in $C_{1}$. Then, we’re done.
  3. (3)
    Our third case is when $C_{1}$ has four vertices, and the two paths $P_{1}$ and $P_{2}$ end at $q$ and $t$, respectively. Let the fourth vertex other than $q,r,$ and $t$ be $u$. In this case, we move $s$ through $P_{2}$ and then to $q$ via $r$, so that $x$ is at $r$, $y$ is at $t$, and $z$ is at the vertex in $P_{2}$ adjacent to $t$. Then, move $s$ around the cycle formed by $P_{1}$, $P_{2}$, and $u$ until $y,z$ are on $u$ and $t$, respectively, with $s$ at $q$. We’re done.
  4. (4)
    Our last case is when $C_{1}$ has three vertices. Here, we may assume without loss of generality that the paths $P_{1}$ and $P_{2}$ end at $q$ and $t$, by permuting our choice of $x$, $y$, and $z$. We’ll show that it’s either possible to pick a new choice of $C_{1}$ which has at least $4$ vertices, or we can instead finish this procedure of moving $\left\{x,y,z\right\}$ to $C_{2}$ while maintaining their adjacency. Let $u^{\prime}$, $v^{\prime}$, and $w^{\prime}$ be an arbitrary path in $C_{2}$; we will move $\left\{x,y,z\right\}$ to $\left\{u^{\prime},v^{\prime},w^{\prime}\right\}$. If cycle $C_{2}$ shares an edge other than $qt$ with $C_{1}$, then we can replace $C_{1}$ with the cycle formed by $P_{1}$, $P_{2}$, and $r$, after which we are done as $C_{1}$ has at least 4 vertices now (and in fact $s$ is on $C_{1}$ as well so we are done with the first step). Otherwise, $C_{2}$ only shares edge $qt$ with $C_{1}$. If at most one of $u^{\prime}$, $v^{\prime}$, and $w^{\prime}$ are in $\left\{q,t\right\}$, then it follows that we can replace the edge $qt$ in $C_{2}$ with the edges $qr,rt$, maintaining the adjacency of $u$, $v$, and $w$ in $C_{2}$. With this new $C_{2}$, we can replace $C_{1}$ with the cycle formed by $P_{1}$, $P_{2}$, and $r$, so that $C_{1}$ and $C_{2}$ still share edges, namely $qr$ and $rt$, and we are done as $C_{1}$ has at least 4 vertices now. Otherwise, two of $u^{\prime}$, $v^{\prime}$, and $w^{\prime}$ coincide with $q$ and $t$; if $C_{2}$ has at least 4 vertices, as letting $u^{\prime}$, $v^{\prime}$, and $w^{\prime}$ be any path in $C_{2}$ will suffice, we can repick $u^{\prime}$, $v^{\prime}$, and $w^{\prime}$ so at most one is in $\left\{q,t\right\}$. So lastly, we have $C_{2}$ has exactly 3 vertices $u^{\prime}$, $v^{\prime}$, and $w^{\prime}$; without loss of generality, say $u^{\prime}=q$, and $v^{\prime}=t$. We will move $\left\{x,y,z\right\}$ to $C_{2}$. Move $s$ through $P_{2}$ and then to $q$ so that $x$ is on $t$ and $z$ is on the vertex adjacent to $t$ in $P_{2}$. Then move $s$ to $w^{\prime}$, then $t$, then along $P_{2}$, so that $x$ is on $w^{\prime}$ and $z$ is on $t$. Then move $s$ along $P_{1}$, then to $w^{\prime}$, $t$, and then $r$ so that $(x,y,z)=(q,t,w^{\prime})$, so $\left\{x,y,z\right\}$ is on $C_{2}$.
If we have not already moved $\{x,y,z\}$ to $C_{2}$, we can now finish this process. Let $P_{1}$ be the shared path of $C_{1}$ and $C_{2}$, $P_{2}$ be the other path in $C_{2}$, and $P_{3}$ be the other path in $C_{1}$. By convention, each path includes the common endpoints. We make use of the fact that $C_{1}$ has at least $4$ vertices, because as argued in case 4, if $C_{1}$ only has 3 vertices then we have already moved $\left\{x,y,z\right\}$ to $C_{2}$. We have multiple cases:
  1. (1)
    If $P_{1}$ has at least three vertices in it, then we finish quite easily: move $s$ around $C_{1}$ until $x,y,z$ are each on $P_{1}$.
  2. (2)
    Otherwise, $P_{1}$ has two vertices, and all vertices of $C_{1}$ are on $P_{3}$, so we can move $s$ around the cycle formed by $P_{2}$ and $P_{3}$ until $x,y,z$ are each in $C_{2}$.
We have shown that it is possible to move $x,y,z$ from $C_{1}$ to $C_{2}$. If $k=3$, following the exact same process, we can move $x,y,z$ from $C_{2}$ to $C_{3}$ while maintaining their adjacency. Thus, regardless of $k$, we have moved $\left\{x,y,z\right\}$ to $C_{k}$. To finish, note that we can move $s$ to $C_{k}$ again using the process which we described before, i.e., the first step. Then, after suitably moving $s$ around $C_{k}$, $\left\{x,y,z\right\}$ will be at $\left\{u,v,w\right\}$.
It can be checked that our algorithm requires $O(n^{2})$ friendly swaps because it requires moving $s$ around cycles $O(1)$ times, where each cycle has at most $n$ vertices and we must move around it at most $n$ times. ∎
4.4. 3-vertex permutation cycles over $\theta$-graphs
In this subsection we will prove the existence of the 3-vertex permutation cycles in $\theta$-graphs defined in Lemma 4.9, restated here.
See 4.9
First we need additional notation. For the graph $H=\theta(i,j,k)$ with $i\leq j\leq k$, let $e_{1}$ and $e_{2}$ denote the endpoints of $H$; $P$ denote the path from $e_{1}$ to $e_{2}$ of length $i$; $Q$ denote the path from $e_{1}$ to $e_{2}$ of length $j$; and $R$ denote the path from $e_{1}$ to $e_{2}$ of length $k$.
For the initial permutation $\sigma$, we will use $p_{1},\cdots,p_{i}$ to denote the $i$ vertices of $\mathsf{Star}_{n}$ located in $\sigma$ on $P$; $q_{1},\cdots,q_{j}$ to denote the vertices of $\mathsf{Star}_{n}$ located on $Q$; and $r_{1},\cdots,r_{k}$ to denote the vertices of $\mathsf{Star}_{n}$ located on $R$; with $p_{1},q_{1}$, and $r_{1}$ being located adjacent to $e_{1}$. Additionally, let $s$ denote the center of $\mathsf{Star}_{n}$, and $t$ denote $\sigma^{-1}(e_{2})$. We may assume without loss of generality that $s$ is on $e_{1}$, since otherwise we can move $s$ to $e_{1}$, perform our 3-cycle, and then revert this move, which adds an extra $O(n)$ moves.
In the remainder of the proof, for convenience, we will overload notation and use $P$, $Q$, and $R$ to refer both to the paths in $H$ as defined previously as well as the permutations over the vertices in $\mathsf{Star}_{n}$ located on those paths. Note that the paths stay the same throughout as we perform friendly swaps, but the permutations change over the course of the operations.
In the following proof, we repeatedly use a sequence of moves where $s$, starting at $e_{1}$, moves along one of the paths $P$, $Q$ or $R$ all the way to $e_{2}$, and then follows another one of the three paths in reverse direction back to $e_{1}$. We call this sequence of moves a type-$AB$ move where $A$ and $B$ are the paths followed in forward and reverse direction respectively.
We break up the proof of the lemma into cases as described in Table 2 below.
[…]
4.9.(i) $i=0,k>1$ $(t,r_{k},r_{k-1})$
4.9.(ii) $i=0,j=k=1$ $(u,q_{1},t)$
4.9.(iii) $i=j=k=1$ $(p_{1},u,t)$
4.9.(iv) $i=j=1$, $k\geq 2$ $(r_{k-1},r_{k},t)$
4.9.(v) $i=1$, $j=2$, $k\geq 3$ $(r_{3},r_{2},r_{1})$
4.9.(vi) $i\geq 1$, $j,k\geq 3$ $(t,r_{k},r_{k-1})$
4.9.(vii) $i=2$, $j=2$, $k<6$ $(r_{k-1},r_{k},t)$
4.9.(viii) $i=2$, $j=2$, $k\geq 6$ $(r_{5},r_{4},r_{3})$
[…]
Before we prove the cases, we state and prove a useful lemma.
Lemma 4.12.
If $i\geq 1$ and $j,k\geq 3$, then $\sigma$ and $\sigma\circ(p_{i}q_{j}r_{k})$ are at distance $O(n)$ in $\mathsf{FS}(\mathsf{Star}_{n},\theta(i,j,k))$.
Proof.
We implement a series of type-$AB$ moves in the following sequence, and keep track of the permutations $P$, $Q$, and $R$.
Move Type Permutation $P$ Permutation $Q$ Permutation $R$ Vertex at $e_{2}$
QPQP $q_{1},p_{1},\ldots,p_{i-1}$ $q_{2},q_{3},\ldots,q_{j},t$ $r_{1},\ldots,r_{k}$ $p_{i}$
QRQR $q_{1},p_{1},\ldots,p_{i-1}$ $q_{3},q_{4},\ldots,q_{j},t,p_{i}$ $q_{2},r_{1},\ldots,r_{k-1}$ $r_{k}$
PQPQ $p_{1},p_{2},\ldots,p_{i-1},r_{k}$ $q_{1},q_{3},\ldots,q_{j},t$ $q_{2},r_{1},r_{2},\ldots,r_{k-1}$ $p_{i}$
RQRQ $p_{1},p_{2},\ldots,p_{i-1},r_{k}$ $q_{2},q_{1},q_{3},\ldots,q_{j}$ $r_{1},r_{2},\ldots,r_{k-1},p_{i}$ $t$
Observe that the four moves above (each of which are performed in $O(n)$ time) are the equivalent of $(q_{1}q_{2})\circ(p_{i}r_{k})$. We call this sequence of moves a $Q$-switch. Define $P$-switches and $R$-switches similarly.
After performing a $Q$-switch, we now perform an $R$-switch. The resulting permutations are as follows: $P$ is $p_{1},p_{2},\ldots,p_{i-1},q_{j}$, $Q$ is $q_{2},q_{1},\ldots,q_{j-1},r_{k}$, and $R$ is $r_{2},r_{1},\ldots,r_{k-1},p_{i}$.
We now repeat these two switches once more in the same order, after which $P$ is $p_{1},p_{2},\ldots,p_{i-1},r_{k}$, $Q$ is $q_{1},q_{2},\ldots,q_{j-1},p_{i}$, and $R$ is $r_{1},r_{2},\ldots,r_{k-1},q_{j}$. We have now achieved $\sigma\circ(p_{i}r_{k}q_{j})$. To reach the desired configuration, we repeat this process once again.
To finish, note that we moved $s$ through each of the paths $P,Q,R$ some constant number of times. Hence, the number of friendly swaps needed to get from the initial configuration to the final configuration is $O(n)$. ∎
We will now prove the cases in Table˜2.
Case 4.9.(i). Recall that we have $i=0$, $j\geq 1$, and $k>1$. Let the permutation $X=x_{1}x_{2}\ldots x_{n-1}$ be the order of the vertices around the path from $r_{1}$ through $r_{k}$ to $e_{2}$ to $q_{j}$ to $q_{1}$. We now follow the below sequence of steps. We display the resulting permutation at the end of each step.
Move Type Resulting configuration
QRQR $x_{n-1},x_{1},x_{2},\ldots,x_{n-2}$
PRPR $x_{k},x_{n-1},x_{1},x_{2},\ldots,x_{k-1},x_{k+1},\ldots,x_{n-2}$
PQPQ $x_{k},x_{n-1},x_{1},x_{2},\ldots,x_{k-2},x_{k+1},\ldots,x_{n-2},x_{k-1}$
RPRP $x_{n-1},x_{1},x_{2},\ldots,x_{k-2},x_{k+1},x_{k},x_{k+2},\ldots,x_{n-2},x_{k-1}$
QPQP $x_{n-1},x_{1},x_{2},\ldots,x_{k-2},x_{k+1},x_{k-1},x_{k},x_{k+2},\ldots,x_{n-2}$
RQRQ $x_{1},x_{2},\ldots,x_{k-2},x_{k+1},x_{k-1},x_{k},x_{k+2},\ldots,x_{n-1}$
We have achieved what we wanted here, and we have done it in $O(n)$ friendly swaps, because we have moved $s$ through each of $P$, $Q$, and $R$ a constant number of times. ∎
Case 4.9.(ii). Recall that we have $i=0,j=k=1$. Here, we use a slightly different setup as all of the other cases. We assume that our vertex $s$ is on the singular vertex in $R$. We’ll denote the vertex located at $e_{1}$ as $u$ and use $t$ and $q_{1}$ to denote the other two vertices as before. Then, moving $s$ to $e_{2}$, back through $P$, and then to its original position results in the action $(uq_{1}t)$. ∎
Case 4.9.(iii). Recall that we have $i=j=k=1$. Like in the previous case, we use a slightly different setup as all of the other cases. We assume that our vertex $s$ is on the singular vertex in $Q$. We’ll denote the vertex at $e_{1}$ as $u$ and $t$ and $p_{1}$ as before. Then the cycle $(stp_{1}u)$ followed by $(st)$ implements the cycle $(p_{1}ut)$. ∎
Case 4.9.(iv). Recall that we have $i=j=1,k\geq 2$. We display a series of moves which proves this case in the table below.
Move Type Permutation $P$ Permutation $Q$ Permutation $R$ Vertex at $e_{2}$
PRPR $t$ $q_{1}$ $p_{1}r_{1}\ldots r_{k-1}$ $r_{k}$
QRQR $t$ $r_{k}$ $q_{1}p_{1}r_{1}\ldots r_{k-2}$ $r_{k-1}$
PQPQ $r_{k-1}$ $t$ $q_{1}p_{1}r_{1}\ldots r_{k-2}$ $r_{k}$
RQRQ $r_{k-1}$ $q_{1}$ $p_{1}r_{1}\ldots r_{k-2}r_{k}$ $t$
RPRP $p_{1}$ $q_{1}$ $r_{1}\ldots r_{k-2}r_{k}t$ $r_{k-1}$
This satisfies the problem, and we have moved $s$ through each of $P$, $Q$, and $R$ a constant number of times, so this operation is doable in $O(n)$ friendly swaps. ∎
Case 4.9.(v). Recall that we have $i=1$, $j=2$, and $k\geq 3$. We’ll fix an orientation of the cycle formed by $P$ and $R$ (which we’ll call $C$) in the direction of $p_{1}$ to $s$ to $r_{1}$. We’ll label $c_{1}=p_{1}$, $c_{2}=r_{1}$, and label the other vertices of $C$ around the cycle $c_{3},\ldots,c_{d}$ in that fashion (but we skip over $s$). We will consider the evolution of the circular permutation of the elements around the cycle except for $s$ starting at $c_{1}=p_{1}$.
First, we will define two operations that involve a sequence of type-$AB$ moves. Operation 1 is defined as follows:
Move Type Permutation $P$ Permutation $Q$ Permutation $R$ Vertex at $e_{2}$
QPQP $q_{1}$ $q_{2},t$ $r_{1},\ldots,r_{k}$ $p_{1}$
RPRP $r_{1}$ $q_{2},t$ $r_{2},\ldots,r_{k},p_{1}$ $q_{1}$
$2\times(QP\circ RP)$ $r_{3}$ $q_{1},q_{2}$ $r_{4},\ldots,r_{k},p_{1},r_{1},r_{2}$ $t$
In the configuration reached through the table, observe that we have changed our circular permutation from $c_{1},c_{2},\ldots,c_{d}$ to $c_{1},c_{2},c_{3},c_{d},c_{4},\ldots,c_{d-1}$. We call this an operation of type 1. In our original configuration, we can similarly make a move of type $RP$, followed by operation 1, and then a move of type $PR$. This makes our permutation go from $c_{1},c_{2},\ldots,c_{d}$ to $c_{2},c_{3},c_{4},c_{1},\ldots,c_{d}$. We call this process Operation 2.
Observe that both operations 1 and 2 take $O(n)$ time. We now focus on the five adjacent elements in our original circular permutation, $c_{d},c_{1},c_{2},c_{3},c_{4}$ in that order (because the positions of the rest of the elements do not change). We display our algorithm in the table below:
Operation Permutation
3 $\times$ Operation 1 $c_{3},c_{d},c_{1},c_{2},c_{4}$
3 $\times$ Operation 2 $c_{3},c_{4},c_{d},c_{1},c_{2}$
2 $\times$ Operation 1 $c_{d},c_{1},c_{3},c_{4},c_{2}$
The action we have performed in the table corresponds to making the move $(c_{2}c_{4}c_{3})=(r_{1}r_{3}r_{2})$ in our circular permutation. To finish, note that we can move $s$ around $C$ until the rest of the elements ($c_{5}$, etc) return to their original positions, after which our new permutation is $\sigma\circ(r_{1}r_{3}r_{2})$. This completes our construction. ∎
Case 4.9.(vi). This is the general case where $i\geq 1$ and $j,k\geq 3$. We first describe the following sequence of moves shown in the table, which we’ll denote as $S_{1}$.
Move Type Permutation $P$ Permutation $Q$ Permutation $R$ Vertex at $e_{2}$
22 $\times$ QRQR $p_{1}\ldots p_{i}$ $q_{3}q_{4}q_{5}\ldots q_{j}tr_{k}$ $q_{2}q_{1}r_{1}r_{2}\ldots r_{k-2}$ $r_{k-1}$
PRPR $p_{2}\ldots p_{i}r_{k-1}$ $q_{3}q_{4}\ldots q_{j}tr_{k}$ $p_{1}q_{2}q_{1}r_{1}r_{2}\ldots r_{k-3}$ $r_{k-2}$
PQPQ $p_{3}\ldots p_{i}r_{k-1}r_{k-2}$ $p_{2}q_{3}q_{4}\ldots q_{j}t$ $p_{1}q_{2}q_{1}r_{1}r_{2}\ldots r_{k-3}$ $r_{k}$
RQRQ $p_{3}\ldots p_{i}r_{k-1}r_{k-2}$ $p_{1}p_{2}q_{3}q_{4}\ldots q_{j}$ $q_{2}q_{1}r_{1}r_{2}\ldots r_{k-3}r_{k}$ $t$
QPQP $p_{1}p_{3}\ldots p_{i}r_{k-1}$ $p_{2}q_{3}q_{4}\ldots q_{j}t$ $q_{2}q_{1}r_{1}r_{2}\ldots r_{k-3}r_{k}$ $r_{k-2}$
Let the moves in Lemma˜4.12 be denoted as $S_{2}$. We note that $S_{1}$ and $S_{2}$ can both be performed in $O(n)$ time.
Observe that after performing $S_{1}\circ S_{2}\circ S_{1}^{-1}$, every vertex is returned to its original position except for the three vertices adjacent to $e_{2}$ before we perform $S_{2}$. These vertices are precisely $r_{k-1}$, $t$, and $r_{k}$. It’s easy to then verify that we have made the move $(tr_{k}r_{k-1})$. Therefore, we are done. ∎
Case 4.9.(vii). By Theorem˜4.3, for $k=2,3,4,5$, we have $\mathsf{FS}(\mathsf{Star}_{k+6},\theta(2,2,k))$ is either connected or has two connected components, where different permutations (always with $s$ at $x$) are in the same connected component if they are of the same parity. Because $\sigma$ and $\sigma\circ(r_{k-1}r_{k}t)$ are the same parity for any $\sigma$, it follows that both are in the same connected component. Furthermore, the diameter of this connected component clearly has to be less than its size, which is at most $11!=O(1)$ vertices. ∎
Case 4.9.(viii). Recall that $i=j=2$ and $k\geq 5$. In this case, we will work with circular permutations again, similarly to case (v). We orient the cycle formed by $P$ and $R$ in the direction from $t$ to $p_{2}$ to $s$ to $r_{1}$ to $r_{k}$. Let the circular permutation of the elements (except for $s$) around the cycle formed by $P$ and $R$ be $C=c_{1},c_{2},\ldots,c_{d}$, where $c_{1}$ is $t$ and $c_{2}$ is $p_{2}$.
We will again define two kinds of operations. Operation 1 is defined as follows:
Move Type Permutation $P$ Perm. $Q$ Perm. $R$ Vertex at $e_{2}$
QPQP $q_{1},p_{1}$ $q_{2},t$ $r_{1},\ldots,r_{k}$ $p_{2}$
22 $\times$ RPRP $r_{2},r_{1}$ $q_{2},t$ $r_{3},\ldots,r_{k},p_{2},p_{1}$ $q_{1}$
22 $\times$ $(QP\circ 2\times RP)$ $r_{6},r_{5}$ $q_{1},q_{2}$ $r_{7},\ldots,r_{k},p_{2},p_{1},r_{1},r_{2},r_{3},r_{4}$ $t$
In operation 1, we see that we have changed our circular permutation from $c_{1},c_{2},\ldots,c_{d}$ to $c_{2},c_{3},c_{4},c_{5},c_{6},c_{7},c_{1},\ldots,c_{d}$.
Operation 2 is defined as follows. We make a move of type $RP$, perform Operation 1, and then make a move of type $PR$. Operation 2 applied to the permutation $c_{1},c_{2},\ldots,c_{d}$ changes it to $c_{1},c_{3},c_{4},c_{5},c_{6},c_{7},c_{8},c_{2},\ldots,c_{d}$.
Observe that both Operation 1 and Operation 2 take $O(n)$ time. We will now focus on the eight adjacent elements in our permutation, $c_{1},c_{2},\ldots,c_{8}$, noting that Operations 1 and 2 do not change the positions of the other elements in our circular permutation. We show an algorithm to rearrange these elements into $c_{1},c_{2},c_{3},c_{4},c_{5},c_{7},c_{8},c_{6}$ in the following table.
Operation Permutation
Operation 1 $c_{2},c_{3},c_{4},c_{5},c_{6},c_{7},c_{1},c_{8}$
5 $\times$ Operation 2 $c_{2},c_{1},c_{8},c_{3},c_{4},c_{5},c_{6},c_{7}$
Operation 1 $c_{1},c_{8},c_{3},c_{4},c_{5},c_{6},c_{2},c_{7}$
Operation 2 $c_{1},c_{3},c_{4},c_{5},c_{6},c_{2},c_{7},c_{8}$
Perform rows 1-4 three more times. $c_{1},c_{6},c_{2},c_{3},c_{4},c_{5},c_{7},c_{8}$
Operation 2 $c_{1},c_{2},c_{3},c_{4},c_{5},c_{7},c_{8},c_{6}$
Similarly to case (v), we see that our algorithm performs the action $(c_{8}c_{7}c_{6})=(r_{5}r_{4}r_{3})$. We can now suitably move $s$ around the cycle formed by $P$ and $R$ so that the unchanged elements in our circular permutation (e.g. $c_{1}$) move back to their original position in $\sigma$. After this, it is clear that we have reached the configuration $\sigma\circ(r_{5}r_{4}r_{3})$, so we are done. ∎
4.5. A lower bound
We now show the existence of a graph $Y$ such that $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is connected but its diameter is $\Omega(n^{3})$. This shows that our diameter bound of Theorem˜1.3 is tight within a factor of $O(n)$.
Definition 4.13.
Let $B_{n}$ be the graph on $n$ vertices with vertex set $\{(i,i+1),i\in[n-1]\}\cup\{(1,n),(1,3)\}$. An example of $B_{n}$ is shown in Figure˜3.
Figure 3. The graph $B_{6}$.
Theorem 4.14.
The diameter of $\mathsf{FS}(\mathsf{Star}_{n},B_{n})$ is $\Omega(n^{3})$.
Proof.
We establish the bound by reducing a circular sorting problem to the problem of finding a path in $\mathsf{FS}(\mathsf{Star}_{n},B_{n})$. A circular permutation is an arrangement of objects in a circle, where the cyclic order of these elements matters but there is no notion of a starting element or ending element, unlike in normal permutations. We define an adjacent swap on a circular permutation to be an action which switches two adjacent elements in our permutation but does not affect the position of any of the other elements. Adin, Alon, and Roichman [1] show the existence of two circular permutations over n1n-1 elements such that the number of adjacent swaps required to transform one into the other is $\Omega(n^{2})$. Let $\sigma$ and $\omega$ be these permutations.
Label the nodes of $\mathsf{Star}_{n}$ by $\{s\}\cup\{x_{1},\cdots,x_{n-1}\}$. We first associate any configuration in $\mathsf{FS}(\mathsf{Star}_{n},B_{n})$ with a circular permutation over $x_{1},\dots,x_{n-1}$ by removing $s$ and considering the ordering of the other elements around the Hamiltonian cycle of $B_{n}$. In particular, a configuration in $\mathsf{FS}(\mathsf{Star}_{n},B_{n})$ can be specified by specifying the location of $s$, the vertex in $\mathsf{Star}_{n}$ adjacent to it in $B_{n}$ in clockwise fashion, and then the remaining circular permutation over $x_{1},\cdots,x_{n-1}$.
We will argue that the distance between any two configurations in $\mathsf{FS}(\mathsf{Star}_{n},B_{n})$, where one is associated with $\sigma$ and the other with $\omega$, is $\Omega(n^{3})$.
Note that any sequence of friendly moves along the main cycle $B_{n}\setminus\{(1,3)\}$ maintains the circular permutation over $x_{1},\cdots,x_{n-1}$. On the other hand, any friendly swap over the edge $(1,3)$ corresponds to an adjacent swap on the corresponding circular permutation (note that this swap is between the vertex at 2 and the vertex at 1 or 3).
Our key claim is that between any two distinct adjacent swaps, we must conduct at least $\Omega(n)$ friendly moves along the main cycle $B_{n}\setminus\{(1,3)\}$. This immediately implies the lower bound of $\Omega(n^{3})$.
To prove the claim, we note that performing an adjacent swap requires $s$ to be located at $1$ or $3$. Suppose that at the end of such a swap, $s$ is located at $3$ and $x_{i}$ and $x_{j}$ are located at $1$ and $2$ respectively. We now make a set of moves along the main cycle $B_{n}\setminus\{(1,3)\}$ before our next swap. We have two cases:
  1. (1)
    If $s$ ends at $1$ after this set of moves, then $s$ must have moved through the path $3,4,\ldots,n$, or otherwise $x_{i}$ and $x_{j}$, respectively, will be at $2$ and $3$, and performing our swap will result in switching the same two elements in our circular permutation.
  2. (2)
    If $s$ ends at $3$ after this set of moves, then $s$ must have moved fully around our cycle. Otherwise, we have made a series of friendly swaps and undone these friendly swaps, so after our moves the configuration is the same as directly after our adjacent swap, and friendly-swapping $s$ and $x_{i}$ will restore our original configuration.
In either case, we conclude that any pair of distinct adjacent swaps requires $\Omega(n)$ intermediate moves to accomplish. ∎
5. General Diameter Bounds Depending on Minimum Degree
In [3], several conditions on the minimum degrees of $X$ and $Y$ are given to guarantee the connectedness of $\mathsf{FS}(X,Y)$. In this section, we will prove two similar theorems, but we will focus on bounding the diameter of our friends-and-strangers graphs rather than proving their connectivity. Our first result is based on a result of Bangachev [3].
Theorem 5.1 ([3]).
Let $X$ and $Y$ be connected graphs on $n$ vertices such that
$\min(\delta(X),\delta(Y))+2\max(\delta(X),\delta(Y))\geq 2n.$
Then, $\mathsf{FS}(X,Y)$ is connected.
We improve on this result by showing that under the same conditions established in this theorem, the diameter of $\mathsf{FS}(X,Y)$ is polynomially bounded.
See 1.5
Proof.
For a graph $G$, we denote the induced subgraph over a subset $S\subset V(G)$ as $G\mid_{S}$. Furthermore, for a vertex $i$ in $G$, define $N[i]$ as the closed neighborhood of $i$, or in other words the union of $\{i\}$ and the set of vertices in $G$ which are adjacent to $i$.
Let $\sigma$ denote a bijection from $V(X)$ onto $V(Y)$, and let $u$ and $v$ be vertices in $Y$, with $\hat{u}\coloneq\sigma(u)$ and $\hat{v}\coloneq\sigma(v)$, such that $\hat{u}$ and $\hat{v}$ are adjacent in $X$. Let $Q$ be $\sigma(N[\sigma^{-1}(u)])$. In [3], it is shown that $\delta(Y|_{Q})>\frac{1}{2}|Q|$.
[…]
For the first fact, we cite another theorem in Bangachev, which says that $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is connected if $\delta(Y)>\frac{n}{2}$. Hence, by Theorem˜1.3 we conclude that the diameter of $\mathsf{FS}(\mathsf{Star}_{|Q|},Y|_{Q}))$ is at most $O(|Q|^{4})\leq O(n^{4})$. Furthermore, $X|_{\sigma(Q)}$ clearly has a subgraph isomorphic to $\mathsf{Star}_{|Q|}$, because $u$ is adjacent to every vertex in this graph. Hence, we conclude that $\mathsf{FS}(\mathsf{Star}_{|Q|},Y|_{Q}))$ is a subgraph of $\mathsf{FS}(X|_{\sigma(Q)},Y|_{Q})$, and the latter graph has diameter $O(n^{4})$. The assertion follows, because this graph is connected and we can therefore clearly switch the positions of $\hat{u}$ and $\hat{v}$ in $O(n^{4})$ moves.
[…]
Our second result provides a much stronger diameter bound than the first result, given a similar but stronger condition on $\delta(X)$ and $\delta(Y)$.
See 1.6
Proof.
Let $\sigma$ denote the bijection from $V(X)$ onto $V(Y)$. Let $p$ and $q$ be vertices in $X$ such that $pq$ is an edge and let $a=\sigma(p)$ and $b=\sigma(q)$. We will prove the existence of a vertex $r$ in $Y$ that satisfies the following properties: (i) $r$ is adjacent to $a$ in $Y$; (ii) $r$ is adjacent to $b$ in $Y$; (iii) $\sigma^{-1}(r)$ is adjacent to $p$ in $X$; and, (iv) $\sigma^{-1}(r)$ is adjacent to $q$ in $X$.
At most $n-1-\delta(Y)$ fail the condition (i); at most $n-1-\delta(Y)$ fail (ii); at most $n-1-\delta(X)$ fail (iii); and at most $n-1-\delta(X)$ fail (iv). In all, at most $4n-4-2\delta(X)-2\delta(Y)<n-2$ vertices fail to satisfy one or more of these conditions. Therefore, there exists at least one vertex other than $a$ and $b$ that satisfies all of these conditions. Call it $r$.
Because $r$ is adjacent to $a$ and $b$ and $\sigma^{-1}(r)$ is adjacent to $p$ and $q$, we can perform the sequence of friendly swaps $(r,p)$, $(p,q)$, $(r,q)$, after which the positions of $p$ and $q$ are swapped. Hence, for any vertices $p$ and $q$ in $X$ for which $pq$ is an edge, it is possible to swap the positions of $p$ and $q$ in $3$ friendly swaps.
Finally, observe that the diameter of $\mathsf{FS}(X,K_{n})$ is at most $\binom{n}{2}=\frac{n(n-1)}{2}$. Hence, if we take a sequence of moves from any configuration in $\mathsf{FS}(X,K_{n})$ to any other configuration, and replace each move with it corresponding sequence of $3$ friendly swaps, we can move from any configuration to any other configuration in $\mathsf{FS}(X,Y)$ in at most $\frac{3n(n-1)}{2}$ friendly swaps. ∎
We remark that Bangachev [3] proves another connectivity result of a similar flavor as the connectivity analog of Theorem˜1.5; see Figure˜1 for a visual comparison of the regimes that these results apply to. We state this result below for completeness.
Theorem 5.2.
[3, Theorem 1.4] Let $X$ and $Y$ be graphs on $n$ vertices. If $\delta(X)>n/2$, $\delta(Y)>n/2$, and $2\min(\delta(X),\delta(Y))+3\max(\delta(X),\delta(Y))\geq 3n$, then $\mathsf{FS}(X,Y)$ is connected.
We leave open the question of finding the diameter analog of this result.
6. Diameter Bounds for Erdős-Rényi Random Graphs
In this section we prove Theorem˜1.4, which we restate below for convenience. See 1.4
As in the work of Alon, Defant and Kravitz [2], our proof relies on the notion of Wilsonian graphs, as defined below. Theorem˜4.3 then shows that if $Y$ is Wilsonian, then $\mathsf{FS}(\mathsf{Star}_{n},Y)$ is connected. We will use this fact in our argument.
Definition 6.1.
We say a graph $G$ is Wilsonian if it is biconnected, non-bipartite, and neither a cycle graph with at least 4 vertices nor isomorphic to $\theta(1,2,2)$.
We will also need the following lemma about biconnectivity in random graphs.
Lemma 6.2.
Let $X\sim\mathcal{G}(n,p)$ with $p\geq 20\log n/n$. Then $X$ is biconnected with probability at least $1-o(n^{-8})$.
Proof.
As having higher $p$ increases the probability that $X$ is biconnected, without loss of generality assume $p=20\log n/n$. We will show that $X$ is biconnected by showing that $X$ does not contain a cut vertex with high probability. We will compute the probability that there exists a vertex $v$ which is a cut vertex. If $v$ is a cut vertex, let $M$ and $N$ be the vertex sets of the connected components of $X\setminus\{v\}$. Assume without loss of generality that $\left\lvert M\right\rvert\leq\left\lvert N\right\rvert$.
We’ll split this into two cases:
  1. (1)
    We first consider the case $\left\lvert M\right\rvert\leq 20$. Fix any nonempty subset $M\subseteq V(X)$ of cardinality at most 20. Then for any vertex in $M$, the number of edges it has with the $n-\left\lvert M\right\rvert\geq n-20$ vertices in $V(X)\setminus M$ is at most 1, namely an edge to $v$. We’ll use Chernoff bounds to bound the probability that any fixed $x\in M$ has at most 1 edge to $V(X)\setminus M$. Note that these events for different $x\in M$ are mutually independent.
    The expected number of edges among these $n-\left\lvert M\right\rvert$ possible edges is
    $\mu=20\log n-\frac{20M\log n}{n}\geq 20\log n-\frac{400\log n}{n},$
    so if we define $\delta=\frac{\mu-1}{\mu}$, then the probability $x\in M$ has at most 1 edge to $V(X)\setminus M$ is at most $e^{-\delta^{2}\mu/2}$. For large enough $n$, we have $\delta^{2}\geq 0.991$, so this value is bounded from above by
    $e^{-0.991\mu/2}<n^{-9.9},$
    where this inequality holds for all sufficiently large $n$.
    To finish this case, note that our analysis does not assume $v$ is fixed; we have shown that for fixed $M$, the probability $X$ has a cut vertex that separates $X$ into two connected components, the smaller of which is $M$, is at most $n^{-9.9\left\lvert M\right\rvert}$. In other words, this $n^{-9.9\left\lvert M\right\rvert}$ bound includes all possible values of $v$. Thus, it only remains to take a union bound over all possibilities for $M$, so the probability of this case occurring is at most
    $\sum_{\left\lvert M\right\rvert=1}^{20}\binom{n}{\left\lvert M\right\rvert}\cdot(n^{-9.9})^{\left\lvert M\right\rvert}=o(n^{-8}).$
  2. (2)
[…]
    $\sum_{r=21}^{n/2-1}\binom{n}{r}\cdot(n-r)\cdot(1-p)^{r(n-1-r)},$
    where the first term represents the number of ways to choose the subgraph $M$, the second term represents the number of ways to choose $v$, and the third term represents the probability that none of the edges between $M$ and $N$ exist. We can bound this from above by
    $\sum_{r=21}^{n/2-1}n^{r+1}(1-p)^{r(n-1-r)}$ $\leq\sum_{r=21}^{n/2-1}\operatorname{exp}((r+1)\log(n)-pr(n-1-r))$
    $\leq\sum_{r=21}^{n/2-1}\operatorname{exp}((r+1)\log(n)-10r\log(n))$
    $\leq\sum_{r=21}^{n/2-1}\operatorname{exp}(-8r\log(n))$
    $=\sum_{r=21}^{n/2-1}n^{-8r}=O(n^{-168}).$
Therefore, the probability that $X$ has a cut vertex is at most $o(n^{-8})$. It follows that the probability that $X$ is biconnected is $1-o(n^{-8})$. ∎
Now we are ready to prove the main result of this section.
Proof of Theorem˜1.4.
Let $X$ and $Y$ be random graphs chosen independently from $\mathcal{G}(n,p)$ and $\mathcal{G}(n,q)$, respectively, over the vertex set $[n]$. Without loss of generality we assume $p\geq q$, so in particular, $p\geq 10\sqrt{\log n/n}$. Let m=(n1)pm=(n-1)p and $\epsilon=\sqrt{10\log n/m}$. Observe that $\epsilon=o(1)$.
We first note that, by Lemma˜6.2, $X$ is connected with probability $1-o(n^{-8})$. Moreover, with high probability the degree of every vertex in $X$ is in $(1-\epsilon)m,(1+\epsilon)m]$. In particular, for a fixed vertex $v\in V(X)$, the expected degree of $v$ in $X$ is $m$, and therefore using Chernoff bounds, the probability that $v$’s degree falls outside of the range $(1-\epsilon)m,(1+\epsilon)m]$ is at most $\operatorname{exp}(-\epsilon^{2}m/2)=n^{-5}$. Taking the union bound over the $n$ vertices, we get that with probability at least $1-n^{-4}$, every vertex in $X$ has degree in $(1-\epsilon)m,(1+\epsilon)m]$ and $X$ is connected. We will condition on this event for the rest of the argument.
[…]
It remains to prove the $1-o(n^{-4})$ lower bound on the probability of occurrence of $\mathcal{E}_{\sigma,a,b}$. Recall that $\hat{a}=\sigma^{-1}(a)$ and $\hat{b}=\sigma^{-1}(b)$ are neighbors. Let $T_{X}$ denote the closed neighborhood of $\hat{a}$ in $X$. Note that $\left\lvert T_{X}\right\rvert\in[(1-\epsilon)m+1,(1+\epsilon)m+1]$. Consider the set of locations of $T_{X}$ under $\sigma$ in $Y$, namely, $T_{Y}=\sigma(T_{X})$. We have $a,b\in T_{Y}$.
[…]
We will now bound the probability that $H_{Y}$ is Wilsonian. We do this by bounding the probability of the three possible ways $H_{Y}$ could not be Wilsonian, where we use $t$ to denote $\left\lvert T_{Y}\right\rvert$.
  • $H_{Y}$ is not biconnected. Note that
    $q\geq\frac{100\log n}{np}>50\frac{\log n}{t}.$
    Then, by Lemma˜6.2, it follows that $H_{Y}$ is not biconnected with probability at most $o(t^{-8})=o(n^{-8/2})$, where we used $t\geq(1-\epsilon)(n-1)p=\Omega(\sqrt{n\log n})$. Therefore, $H_{Y}$ is biconnected with probability $1-o(n^{-4})$.
  • $H_{Y}$ is bipartite. We can bound the probability of this event by summing over all possible bipartitions the probability that the bipartition arises:
    $\sum_{j=0}^{t}{t\choose j}(1-q)^{{j\choose 2}+{{t-j}\choose 2}}\leq 2^{t}(1-q)^{t^{2}/4}<2^{t}e^{-qt^{2}/4}<\operatorname{exp}(t-t\log n)=o(n^{-4}),$
    where we used the fact that $qt\geq(1-\epsilon)(n-1)pq>4\log n$ and $t>5$.
  • $H_{Y}$ is a cycle. For any fixed vertex in $H_{Y}$, its expected degree is (t1)q(t-1)q. We can bound the probability that its degree is at most 2<(t1)q/42<(t-1)q/4 using the Chernoff bound by $\operatorname{exp}(-(t-1)q/8)=o(n^{-4})$. The probability that $H_{Y}$ is a cycle is no more than the probability that one of its vertices has degree at most $2$.
Taking the union bound over the three bad events and applying Definition˜6.1, we get that the probability that $H_{Y}$ is not Wilsonian is at most $o(n^{-4})$, and so we are done. ∎
7. Open Questions
In this section, we state several open questions which we hope to see addressed in future work. The first two questions were proposed by Jeong [7]. This first question is related to Jeong’s theorem that showed that $\mathsf{FS}(X,Y)$ can have connected components of diameter $e^{\Omega(n)}$.
Question 7.1 ([7]).
Let $X$ and $Y$ be graphs on $n$ vertices. Is it true that the maximum possible diameter of a connected component of $\mathsf{FS}(X,Y)$ is $e^{O(n)}$?
Jeong’s friends-and-strangers graph construction for the equality case in the above question has many connected components. This raises a natural followup question.
Question 7.2 ([7]).
Is it true that if $\mathsf{FS}(X,Y)$ is connected, then its diameter is polynomially bounded in terms of $n$?
We’re also interested in improving our diameter bound on $\mathsf{FS}(\mathsf{Star}_{n},Y)$. In this paper, we showed that the bound is $O(n^{4})$, but we have only been able to show that it is bounded from below by $\Omega(n^{3})$. This raises the following natural question.
Question 7.3.
Do there exist graphs $Y$ for which $\mathsf{FS}(\mathsf{Star}_{n},Y)$ has diameter $\Omega(n^{4})$? If not, what is the maximum possible diameter $\mathsf{FS}(\mathsf{Star}_{n},Y)$ can have?
We can ask a similar but more general question about trees, as follows.
Question 7.4.
Let $T$ be a tree and $Y$ be an arbitrary graph, both with $n$ vertices. Is the diameter of any connected component of $\mathsf{FS}(T,Y)$ polynomially bounded in $n$?
We would also like to extend our results in Section˜5, as shown in Figure˜1. We showed that the diameter of $\mathsf{FS}(X,Y)$ is polynomial in $n$ when $\min(\delta(X),\delta(Y))+2\max(\delta(X),\delta(Y))\geq 2n$. On the other hand, Bangachev [3] showed that $\mathsf{FS}(X,Y)$ is connected when $\delta(X),\delta(Y)>\frac{n}{2}$ and $2\min(\delta(X),\delta(Y))+3\max(\delta(X),\delta(Y))\geq 3n$. This result covers many graphs which are not shown to have polynomial diameter in Theorem˜1.5. We ask whether a counterpart of Theorem˜1.5 extends to this regime.
Question 7.5.
Is it true that when $\delta(X),\delta(Y)>\frac{n}{2}$ and $2\min(\delta(X),\delta(Y))+3\max(\delta(X),\delta(Y))\geq 3n$, $\operatorname{diam}(\mathsf{FS}(X,Y))$ is polynomially bounded in $n$?
Bangachev [3] also conjectured that the condition $\delta(X),\delta(Y)>\frac{n}{2}$ is not necessary for connectivity. Therefore, we ask the following natural question.
Question 7.6.
Is it true that when $2\min(\delta(X),\delta(Y))+3\max(\delta(X),\delta(Y))\geq 3n$, $\operatorname{diam}(\mathsf{FS}(X,Y))$ is polynomially bounded in $n$?
In Section˜6, we showed that when $X$ and $Y$ are Erdős-Rényi random graphs satisfying certain conditions, the distance between any two permutations in $\mathsf{FS}(X,Y)$ is $O(n^{6})$ with high probability. We’re curious to know if in the same regime, one can establish polynomial diameter bounds with high probability.
Question 7.7.
Let $p$ and $q$ be probabilities and $c$ a constant for which $pq\geq\frac{c\log{n}}{n}$. Is it true with high probability that $\mathsf{FS}(\mathcal{G}(n,p),\mathcal{G}(n,q))$ has connected components with $\operatorname{poly}(n)$ diameter?
There are many other questions about friends-and-strangers graphs which are still open; we refer the interested reader to Jeong [7] for a list with additional open questions.
Acknowledgments
Rupert Li was partially supported by a Hertz Fellowship and a PD Soros Fellowship. We are grateful to the PRIMES-USA program and its organizers, Dr. Tanya Khovanova, Prof. Pavel Etingof, and Dr. Slava Gerovitch, for providing us the opportunity to conduct this research. Furthermore, we thank Dr. Khovanova for proofreading our work and providing helpful feedback.